# A Unified Theory of Sparse Dictionary Learning in Mechanistic Interpretability: Piecewise Biconvexity and Spurious Minima

### v3

@@ 5.2 Results on CLIP Embeddings of ImageNet
Table 2: Feature recovery on the CLIP embeddings.
[…]
TopK SAE 0.00% 0.517 64.0
   + Feature Anchoring 24.13% 0.851 64.0
BatchTopK SAE 0.00% 0.508 64.0
   + Feature Anchoring 24.13% 0.847 64.0
Matryoshka SAE 0.00% 0.683 128.0
   + Feature Anchoring 24.45% 0.858 128.0
@@ 6 Conclusion
We provide full scripts of the Linear Representation Bench and feature anchoring in this GitHub repository: The_Linear_Representation_Bench.
@@ Appendix B Notations
We summarize the key notations used throughout this paper.
@@
$\mathbf{S}$ Similarity matrix: $|W_{p}^{\text{learned}\top}W_{p}^{\text{true}}|$
@@ Appendix C The Linear Representation Bench
The Linear Representation Bench is a synthetic benchmark designed to precisely instantiate the Linear Representation Hypothesis (Assumption 2.3) and Representation Assumptions (Assumption 2.4) with fully accessible ground-truth features (See Figure 8 for a visualization of the benchmark). This enables rigorous evaluation of SDL methods under controlled conditions where feature recovery can be directly measured.
@@ C.1 Ground-Truth Feature Matrix Generation
We construct the feature matrix $\mathbf{W}_{p}\in\mathbb{R}^{n_{p}\times n}$ to satisfy the unit-norm and bounded interference conditions (Assumption 2.4).
@@ C.3 Data Synthesis
By construction, this dataset exactly satisfies Assumptions 2.3 and 2.4, enabling evaluation of SDL methods with complete knowledge of ground-truth features.
@@ D.1 Mechanistic Interpretability
Interpretability is crucial for deploying AI in high-stakes domains such as medical diagnosis and financial modeling, where understanding model decisions is essential for safety and trust (Simon and Zou, 2024; Abdulaal et al., 2024). Traditional approaches include interpretability-by-design methods like Concept Bottleneck Models (Koh et al., 2020) and decision trees (Mienye and Jere, 2024), and post-hoc explanation methods like GradCAM (Selvaraju et al., 2019) and SHAP (Lundberg and Lee, 2017). Mechanistic interpretability (Sharkey et al., 2025; Bereska and Gavves, 2024) aims to reverse-engineer neural networks by understanding their internal computational mechanisms. SAEs (Shu et al., 2025) and related dictionary learning methods (Tang et al., 2025a; Dunefsky et al., 2024) decompose neural activations into sparse, interpretable features. Circuit analysis (Olah et al., 2020; Olsson et al., 2022) investigates how these features compose into computational algorithms.
@@ D.2 Sparse Dictionary Learning
Sparse dictionary learning has a rich history predating its application to mechanistic interpretability. K-SVD (Aharon et al., 2006) established foundational methods for learning overcomplete dictionaries, while theoretical work in compressed sensing (Donoho, 2006) characterized recovery conditions, with Spielman et al. (2012) providing polynomial-time algorithms for exact reconstruction under sparsity assumptions. Safran and Shamir (2018) demonstrated that spurious local minima are common even in simple two-layer ReLU networks, highlighting optimization challenges that persist in modern applications. Recent work has adapted these principles to mechanistic interpretability. SAEs (Cunningham et al., 2023) apply dictionary learning to language model activations. Various variants are further developed, including transcoders (Dunefsky et al., 2024), crosscoders (Gao et al., 2024), Matryoshka SAEs (Bussmann et al., 2025), and hybrid approaches like Language-Grounded Sparse Encoders (Tang et al., 2025a). These methods have found applications beyond language models, including protein structure analysis (Simon and Zou, 2024; Gujral et al., 2025), medical imaging (Abdulaal et al., 2024), model evaluation (Tang et al., 2025b), board game analysis (Karvonen et al., 2024), and fMRI data analysis Mao et al. (2025).
@@ Appendix E Limitations
  • Assumption Violations. Our analysis relies on Assumptions 2.3 and 2.4, which may not hold perfectly in real-world neural networks. In particular, there exists features that are not one-dimentional linear (Engels et al., 2025) .
  • Extreme Sparsity. Several key results, including Theorem 3.1, rely on the extreme sparsity assumption ($S\to 1$). The bounds may degrade substantially for moderate sparsity levels commonly observed in practice.
@@ Appendix G Ablation Studies
To examine how feature anchoring performs under different interference levels, we fix $n=1000$ features in $n_{p}=n_{r}=768$ dimensions and vary the maximum interference $M\in\{0.05,0.1,0.2,0.5\}$ during ground-truth feature matrix generation. Recall from Assumption 2.4 that $M:=\max_{i\neq j}\langle W_{p}[:,i],W_{p}[:,j]\rangle$ quantifies the maximum dot product between distinct feature directions, characterizing the degree of feature overlap in the representation space.
@@
We investigate how feature sparsity $S$ affects SDL performance by varying $S\in\{0.005,0.01,0.05,0.1\}$ where $S$ denotes the probability that each feature is inactive (Assumption 2.3). Lower $S$ values correspond to denser activation patterns, while higher $S$ approaches the extreme sparsity regime analyzed in our theoretical results.
@@
These comprehensive ablation studies validate our theoretical framework across diverse conditions. The consistent failure of ReLU and JumpReLU SAEs without sufficient sparsity constraints corroborates Theorem 3.4: the underdetermined solution space admits infinitely many configurations achieving low reconstruction loss without recovering interpretable features. Feature anchoring addresses this by constraining encoder-decoder pairs, reducing degrees of freedom and steering optimization away from spurious partial minima (Theorem 3.7). The varying effectiveness across interference levels, sparsity regimes, and activation functions reflects the fundamental trade-off in Assumption 2.4: more severe feature compression (higher $M$), denser activations (lower $S$), or insufficient activation sparsity (higher $k$) create more complex optimization landscapes where even anchored methods struggle to disentangle all features completely.
@@ Appendix H Qualitative Examples
Beyond quantitative metrics, we provide qualitative evidence demonstrating how feature anchoring improves feature monosemanticity. Figure 9 shows features learned by Matryoshka SAE with feature anchoring on CLIP embeddings of ImageNet, while Figure 10 shows features from the same architecture trained without feature anchoring.
@@
Figure 9: Features learned with feature anchoring exhibit monosemanticity. Each row shows the top-activating images for a single feature from Matryoshka SAE trained on CLIP embeddings with feature anchoring.
@@ Appendix I Proofs of the Theorems
We establish how the loss decomposes when $S\to 1$. Let $W_{r}=[w_{r}^{1},w_{r}^{2},\ldots,w_{r}^{n}]$ and $W_{p}=[w_{p}^{1},w_{p}^{2},\ldots,w_{p}^{n}]$ denote the column representations of the reconstruction and projection matrices.
Under extreme sparsity ($S\to 1$), we only need to consider situations where at most one component $x_{d}(s)$ is active for any given input. The expectation decomposes over sparsity levels:
$\mathcal{L}_{\text{SDL}}$ $=\mathbb{E}_{s}\left\|W_{r}\mathbf{x}(s)-W_{D}\sigma(W_{E}W_{p}\mathbf{x}(s))\right\|^{2}$ (32)
$=\sum_{m=0}^{n}\Pr(\|\mathbf{x}(s)\|_{0}=m)\cdot\mathbb{E}[\|\mathbf{x}_{r}(s)-W_{D}\sigma(W_{E}\mathbf{x}_{p}(s))\|^{2}\mid\|\mathbf{x}(s)\|_{0}=m]$ (33)
When $\|\mathbf{x}(s)\|_{0}=0$, both $\mathbf{x}_{r}(s)=0$ and the reconstruction $W_{D}\sigma(W_{E}\mathbf{x}_{p}(s))=0$, contributing zero to the loss.
When $\|\mathbf{x}(s)\|_{0}=1$, exactly one feature is active: $\mathbf{x}(s)=x_{d}(s)\mathbf{e}_{d}$ for some $d\in[n]$. In this case:
$\mathbf{x}_{r}(s)$ $=W_{r}\mathbf{x}(s)=x_{d}(s)w_{r}^{d}$ (34)
$\mathbf{x}_{p}(s)$ $=W_{p}\mathbf{x}(s)=x_{d}(s)w_{p}^{d}$ (35)
Therefore:
$\sum_{d=1}^{n}\Pr(\mathbf{x}(s)=x_{d}(s)\mathbf{e}_{d})\cdot\mathbb{E}[\|\mathbf{x}_{r}(s)-W_{D}\sigma(W_{E}\mathbf{x}_{p}(s))\|^{2}\mid\mathbf{x}(s)=x_{d}(s)\mathbf{e}_{d}]$ (36)
$=\sum_{d=1}^{n}\Pr(\mathbf{x}(s)=x_{d}(s)\mathbf{e}_{d})\cdot\mathbb{E}_{x_{d}(s)}\left[x_{d}(s)^{2}\left\|w_{r}^{d}-W_{D}\sigma(W_{E}w_{p}^{d})\right\|^{2}\mid x_{d}(s)>0\right]$ (37)
$=\sum_{d=1}^{n}M_{d}\left\|w_{r}^{d}-W_{D}\sigma(W_{E}w_{p}^{d})\right\|^{2}$ (38)
where $M_{d}:=\Pr(\mathbf{x}(s)=x_{d}(s)\mathbf{e}_{d})\cdot\mathbb{E}[x_{d}(s)^{2}|x_{d}(s)>0]$.
When $\|\mathbf{x}(s)\|_{0}\geq 2$, by Assumption 2.4 (extreme sparsity), $\sum_{m=2}^{n}\Pr(\|\mathbf{x}(s)\|_{0}=m)\leq(1-S)^{2}$. Since the reconstruction error is bounded by a constant $C$ (due to bounded norms), this contribution is $O((1-S)^{2})$.
Combining all terms:
$\mathcal{L}_{\text{SDL}}=\sum_{d=1}^{n}M_{d}\left\|w_{r}^{d}-W_{D}\sigma(W_{E}w_{p}^{d})\right\|^{2}+O((1-S)^{2})$ (39)
@@
Notation. Recall that $\mathcal{A}(d)\subseteq[n_{q}]$ denotes the set of neurons activated when feature $d$ is active in isolation. Formally, for $W_{E}\in\Omega_{\mathcal{A}}$:
$\mathcal{A}(d)=\{i\in[n_{q}]:\langle w_{E}^{i},w_{p}^{d}\rangle>c\}$ (40)
where $c\geq 0$ is the activation threshold.
Preliminary observation. Within $\Omega_{\mathcal{A}}$, the activation $z_{d}(W_{E}):=\sigma(W_{E}w_{p}^{d})$ satisfies:
$z_{d}(W_{E})^{i}=\begin{cases}\langle w_{E}^{i},w_{p}^{d}\rangle&\text{if }i\in\mathcal{A}(d)\\ 0&\text{if }i\notin\mathcal{A}(d)\end{cases}$ (41)
Therefore, $z_{d}(W_{E})$ is a linear function of $W_{E}$ within $\Omega_{\mathcal{A}}$.
Per-feature loss decomposition. Define the per-feature reconstruction loss:
$f_{d}(W_{D},W_{E}):=\|w_{r}^{d}-W_{D}\sigma(W_{E}w_{p}^{d})\|^{2}$ (42)
The approximate SDL loss is a non-negative weighted sum:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})=\sum_{d=1}^{n}M_{d}f_{d}(W_{D},W_{E})$ (43)
where $M_{d}>0$ for all $d\in[n]$.
We prove biconvexity by establishing that each $f_{d}$ is biconvex, from which the result follows since positive linear combinations preserve biconvexity.
Part 1: $f_{d}$ is convex in $W_{D}$ for fixed $W_{E}\in\Omega_{\mathcal{A}}$.
Fix $W_{E}\in\Omega_{\mathcal{A}}$ and let $z_{d}:=z_{d}(W_{E})$ be the fixed activation vector. Then:
$f_{d}(W_{D})=\|w_{r}^{d}-W_{D}z_{d}\|^{2}$ (44)
Gradient. The gradient with respect to $W_{D}$ is:
$\nabla_{W_{D}}f_{d}=-2(w_{r}^{d}-W_{D}z_{d})z_{d}^{\top}$ (45)
Hessian. The Hessian (treating $W_{D}$ as vectorized) is:
$\nabla^{2}_{W_{D}}f_{d}=2I_{n_{r}}\otimes(z_{d}z_{d}^{\top})$ (46)
Since $z_{d}z_{d}^{\top}\succeq 0$, the Kronecker product with $I_{n_{r}}$ is PSD, confirming $f_{d}$ is convex in $W_{D}$.
Part 2: $f_{d}$ is convex in $W_{E}$ for fixed $W_{D}$, within $\Omega_{\mathcal{A}}$.
Fix $W_{D}\in\mathbb{R}^{n_{r}\times n_{q}}$. Within $\Omega_{\mathcal{A}}$, using the linearity of $z_{d}(W_{E})$:
$W_{D}z_{d}(W_{E})=\sum_{i\in\mathcal{A}(d)}\langle w_{E}^{i},w_{p}^{d}\rangle w_{D}^{i}=:u_{d}(W_{E})$ (47)
where $u_{d}(W_{E})$ is linear (hence affine) in $W_{E}$. Thus:
$f_{d}(W_{E})=\|w_{r}^{d}-u_{d}(W_{E})\|^{2}$ (48)
Since the squared norm of an affine function is convex, $f_{d}$ is convex in $W_{E}$.
Verification via Hessian. For each encoder row $w_{E}^{i}$ with $i\in\mathcal{A}(d)$, the gradient is:
$\nabla_{w_{E}^{i}}f_{d}=-2(w_{r}^{d}-u_{d}(W_{E}))^{\top}w_{D}^{i}\cdot w_{p}^{d}$ (49)
The Hessian is:
$\nabla^{2}_{w_{E}^{i}}f_{d}=2\|w_{D}^{i}\|^{2}(w_{p}^{d}(w_{p}^{d})^{\top})$ (50)
Since $w_{p}^{d}(w_{p}^{d})^{\top}\succeq 0$ and $\|w_{D}^{i}\|^{2}\geq 0$, we have $\nabla^{2}_{w_{E}^{i}}f_{d}\succeq 0$. For $i\notin\mathcal{A}(d)$, neuron $i$ does not affect $f_{d}$, so $\nabla^{2}_{w_{E}^{i}}f_{d}=0$. The full Hessian $\nabla^{2}_{W_{E}}f_{d}$ is block-diagonal with PSD blocks, hence $\nabla^{2}_{W_{E}}f_{d}\succeq 0$.
Conclusion. Since each $f_{d}$ is biconvex and $M_{d}>0$ for all $d$, the weighted sum:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})=\sum_{d=1}^{n}M_{d}f_{d}(W_{D},W_{E})$ (51)
is biconvex over $\mathbb{R}^{n_{r}\times n_{q}}\times\Omega_{\mathcal{A}}$, as positive linear combinations preserve biconvexity. ∎
@@
We analyze the approximate loss $\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})$ by examining the reconstruction error for each ground-truth feature when active in isolation.
Consider feature $d\in[n]$ active in isolation. The encoder output is $W_{E}^{*}w_{p}^{d}=W_{p}^{\top}w_{p}^{d}$ with $k$-th component:
$(W_{E}^{*}w_{p}^{d})_{k}=\langle w_{p}^{k},w_{p}^{d}\rangle$ (52)
Under Assumption 2.4, the unit-norm condition gives $(W_{E}^{*}w_{p}^{d})_{d}=1$. For $k\neq d$, the bounded interference condition implies $\langle w_{p}^{k},w_{p}^{d}\rangle\leq M$, where $M$ is the maximum interference. Combined with the Cauchy-Schwarz inequality, we have:
$\langle w_{p}^{k},w_{p}^{d}\rangle\in[-1,M]\quad\forall k\neq d$ (53)
Define the set of activated neurons for feature $d$:
$\mathcal{A}_{d}=\{k\in[n]:\sigma(W_{E}^{*}w_{p}^{d})_{k}>0\}$ (54)
Since $\sigma(z)_{i}\in\{0,z_{i}\}$ and typical activation functions satisfy $\sigma(z)_{i}>0\iff z_{i}>c$ for some threshold $c\geq 0$, we have:
$\mathcal{A}_{d}=\{k\in[n]:\langle w_{p}^{k},w_{p}^{d}\rangle>c\}$ (55)
Note that $d\in\mathcal{A}_{d}$ since $(W_{E}^{*}w_{p}^{d})_{d}=1>c\geq 0$. For $k\neq d$, neuron $k$ activates only if $\langle w_{p}^{k},w_{p}^{d}\rangle>c\geq 0$, which requires positive interference. Therefore:
$\mathcal{A}_{d}\setminus\{d\}\subseteq\{k\in[n]:0<\langle w_{p}^{k},w_{p}^{d}\rangle\leq M\}$ (56)
Let $K_{d}=|\mathcal{A}_{d}|$ denote the number of activated neurons.
The activation output satisfies $\sigma(W_{E}^{*}w_{p}^{d})_{k}=\langle w_{p}^{k},w_{p}^{d}\rangle$ for $k\in\mathcal{A}_{d}$ and zero otherwise, since $\sigma(z)_{i}\in\{0,z_{i}\}$. The reconstruction is therefore $W_{D}^{*}\sigma(W_{E}^{*}w_{p}^{d})=\sum_{k\in\mathcal{A}_{d}}\langle w_{p}^{k},w_{p}^{d}\rangle w_{r}^{k}$. The reconstruction error for feature $d$ is
$w_{r}^{d}-W_{D}^{*}\sigma(W_{E}^{*}w_{p}^{d})$ $=w_{r}^{d}-\sum_{k\in\mathcal{A}_{d}}\langle w_{p}^{k},w_{p}^{d}\rangle w_{r}^{k}$ (57)
$=w_{r}^{d}-w_{r}^{d}-\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}\langle w_{p}^{k},w_{p}^{d}\rangle w_{r}^{k}$ (58)
$=-\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}\langle w_{p}^{k},w_{p}^{d}\rangle w_{r}^{k}$ (59)
Under the unit-norm assumption, $\|w_{r}^{k}\|=1$ for all $k$. For $k\in\mathcal{A}_{d}\setminus\{d\}$, we have $\langle w_{p}^{k},w_{p}^{d}\rangle>0$ (since negative interference does not activate) and $\langle w_{p}^{k},w_{p}^{d}\rangle\leq M$ by the bounded interference condition. Therefore, using the triangle inequality:
$\left\|w_{r}^{d}-W_{D}^{*}\sigma(W_{E}^{*}w_{p}^{d})\right\|$ $\leq\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}|\langle w_{p}^{k},w_{p}^{d}\rangle|\cdot\|w_{r}^{k}\|$ (60)
$=\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}\langle w_{p}^{k},w_{p}^{d}\rangle\cdot 1\quad\text{(positive terms)}$ (61)
$\leq\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}M$ (62)
$=M(K_{d}-1)$ (63)
Squaring both sides: $\|w_{r}^{d}-W_{D}^{*}\sigma(W_{E}^{*}w_{p}^{d})\|^{2}\leq M^{2}(K_{d}-1)^{2}$. The approximate loss is
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=\sum_{d=1}^{n}M_{d}\cdot M^{2}(K_{d}-1)^{2}\leq M^{2}\sum_{d=1}^{n}M_{d}(K_{d}-1)^{2}$ (64)
Since $K_{d}\geq 1$ for all $d$ and typically $K_{d}=O(1)$ for most features (as the activation function provides sparsity), we have $(K_{d}-1)^{2}=O(1)$. Under the feature independence and stability assumptions (Assumption 2.4), $\sum_{d=1}^{n}M_{d}=\sum_{d=1}^{n}\Pr(\mathbf{x}(s)=x_{d}(s)e_{d})\cdot\mathbb{E}[x_{d}(s)^{2}|x_{d}(s)>0]=O(1-S)$, since the probability that exactly one feature is active is (1S)(1-S) under extreme sparsity. Therefore:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=O(M^{2}(1-S))$ (65)
This completes the proof. ∎

### v5

@@ 5.2 Results on CLIP Embeddings of ImageNet
Table 2: Feature recovery on CLIP embeddings. As ground-truth features are not explicitly available in this setting, we use subpopulation mean embeddings as a proxy for the underlying feature directions. *: 0% recovery resulted from the same issue as Talble 1.
Method $\mathbf{M_{GT}}$ $\uparrow$ $\mathbf{M_{IP}}$ $\uparrow$
TopK SAE 0.00%* 0.517
   + Feature Anchoring 24.13% 0.851
BatchTopK SAE 0.00%* 0.508
   + Feature Anchoring 24.13% 0.847
Matryoshka SAE 0.00%* 0.683
   + Feature Anchoring 24.45% 0.858
@@ 6 Conclusion
We will provide full scripts of the Linear Representation Bench and feature anchoring upon acceptance.
Declaration of LLM Usage
We admit the usage of Claude Sonnet-4.6 for coding and polishing the manuscript.
@@ Appendix C The Linear Representation Bench
The Linear Representation Bench is a synthetic benchmark designed to precisely instantiate the Representation Assumptions (Assumption 2.1) with fully accessible ground-truth features (See Figure 9 for a visualization of the benchmark). This enables rigorous evaluation of SDL methods under controlled conditions where feature recovery can be directly measured.
@@ C.1 Ground-Truth Feature Matrix Generation
We construct the feature matrix $\mathbf{W}_{p}\in\mathbb{R}^{n_{p}\times n}$ to satisfy the representation assumptions (Assumption 2.1).
@@ C.3 Data Synthesis
By construction, this dataset exactly satisfies Assumptions 2.1, enabling evaluation of SDL methods with complete knowledge of ground-truth features.
@@ D.1 Mechanistic Interpretability
Interpretability is crucial for deploying AI in high-stakes domains such as medical diagnosis and financial modeling, where understanding model decisions is essential for safety and trust (Simon and Zou, 2024; Abdulaal et al., 2024; Zhao et al., 2026). Traditional approaches include interpretability-by-design methods like Concept Bottleneck Models (Koh et al., 2020) and decision trees (Mienye and Jere, 2024). Concept Activation Vectors (Kim et al., 2018) extend this by identifying human-defined concept directions in neural network representation spaces. Post-hoc explanation methods like GradCAM (Selvaraju et al., 2019) and SHAP (Lundberg and Lee, 2017) provide local explanations without modifying the model architecture. Mechanistic interpretability (Sharkey et al., 2025; Bereska and Gavves, 2024) aims to reverse-engineer neural networks by understanding their internal computational mechanisms. SAEs (Shu et al., 2025) and related dictionary learning methods (Tang et al., 2025a; Dunefsky et al., 2024) decompose neural activations into sparse, interpretable features. Circuit analysis (Olah et al., 2020; Olsson et al., 2022) investigates how these features compose into algorithms.
@@ D.2 Sparse Dictionary Learning
Sparse dictionary learning has a rich history predating its application to mechanistic interpretability. K-SVD (Aharon et al., 2006) established foundational methods for learning overcomplete dictionaries, while theoretical work in compressed sensing (Donoho, 2006) characterized recovery conditions, with Spielman et al. (2012) providing polynomial-time algorithms for exact reconstruction under sparsity assumptions. Safran and Shamir (2018) demonstrated that spurious local minima are common even in simple two-layer ReLU networks, highlighting optimization challenges that persist in modern applications. Recent work has adapted these principles to mechanistic interpretability. SAEs (Cunningham et al., 2023) apply dictionary learning to language model activations, with LLMs widely utilized to analyze discovered features (Luo et al., 2024a; ; Tang and Dong, 2024; Zou et al., 2026). Various variants are further developed, including transcoders (Dunefsky et al., 2024), crosscoders (Gao et al., 2024), Matryoshka SAEs (Bussmann et al., 2025), and hybrid approaches like Language-Grounded Sparse Encoders (Tang et al., 2025a). These methods have found applications beyond language models, including protein structure analysis (Simon and Zou, 2024; Gujral et al., 2025), medical imaging (Abdulaal et al., 2024), model evaluation (Tang et al., 2025b), prompt engineering (Saini et al., 2026), board game analysis (Karvonen et al., 2024), and fMRI data analysis Mao et al. (2025).
@@ Appendix E Limitations
  • Assumption Violations. Our analysis relies on the Representation Assumptions (Assumption 2.1), which may not hold perfectly in real-world neural networks. In particular, there exists features that are not one-dimentional linear (Engels et al., 2025) .
  • Extreme Sparsity. Several key results, including Theorem 3.1, rely on the extreme sparsity assumption ($1-S<\frac{1}{n}$). The bounds may degrade substantially for moderate sparsity levels commonly observed in practice.
@@ Appendix G Ablation Studies
To examine how feature anchoring performs under different interference levels, we fix $n=1000$ features in $n_{p}=n_{r}=768$ dimensions and vary the maximum interference $M\in\{0.05,0.1,0.2,0.5\}$ during ground-truth feature matrix generation. Recall that $M:=\max_{i\neq j}\langle W_{p}[:,i],W_{p}[:,j]\rangle$ quantifies the maximum dot product between distinct feature directions, characterizing the degree of feature overlap in the representation space.
@@
We investigate how feature sparsity $S$ affects SDL performance by varying $S\in\{0.005,0.01,0.05,0.1\}$ where $S$ denotes the probability that each feature is inactive. Lower $S$ values correspond to denser activation patterns, while higher $S$ approaches the extreme sparsity regime analyzed in our theoretical results.
@@
These comprehensive ablation studies validate our theoretical framework across diverse conditions. The consistent failure of ReLU and JumpReLU SAEs without sufficient sparsity constraints corroborates Theorem 3.4: the underdetermined solution space admits infinitely many configurations achieving low reconstruction loss without recovering interpretable features. Feature anchoring addresses this by constraining encoder-decoder pairs, reducing degrees of freedom and steering optimization away from spurious partial minima (Theorem 3.7). The varying effectiveness across interference levels, sparsity regimes, and activation functions reflects the fundamental trade-off in our assumptions: more severe feature compression (higher $M$), denser activations (lower $S$), or insufficient activation sparsity (higher $k$) create more complex optimization landscapes where even anchored methods struggle to disentangle all features completely.
@@ Appendix H Qualitative Examples
Beyond quantitative metrics, we provide qualitative evidence demonstrating how feature anchoring improves feature monosemanticity. Figure 10 shows features learned by Matryoshka SAE with feature anchoring on CLIP, while Figure 11 shows features from the same architecture trained without feature anchoring.
@@
Figure 10: Features learned with feature anchoring exhibit monosemanticity. Each row shows the top-activating images for a single feature from Matryoshka SAE trained on CLIP with feature anchoring.
@@ Appendix I Proofs of the Theorems
Let $C=\sup_{s\sim\mathcal{D}}\|\mathbf{x}_{r}(s)-W_{D}\sigma(W_{E}\mathbf{x}_{p}(s))\|_{2}^{2}<\infty$ denote the uniform upper bound on the squared reconstruction error. The loss decomposes over the number of active features:
$\mathcal{L}_{\text{SDL}}$ $=\sum_{m=0}^{n}\Pr(\|\mathbf{x}(s)\|_{0}=m)\cdot\mathbb{E}\bigl[\|\mathbf{x}_{r}(s)-W_{D}\sigma(W_{E}\mathbf{x}_{p}(s))\|^{2}\mid\|\mathbf{x}(s)\|_{0}=m\bigr].$ (33)
The $m=0$ term contributes zero. For $m=1$, exactly one feature $d$ is active, giving $\mathbf{x}_{r}(s)=x_{d}(s)w_{r}^{d}$ and $\mathbf{x}_{p}(s)=x_{d}(s)w_{p}^{d}$, so the total contribution equals:
$\sum_{d=1}^{n}M_{d}\left\|w_{r}^{d}-W_{D}\sigma(W_{E}w_{p}^{d})\right\|^{2}=\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E}),$ (34)
where $M_{d}:=\Pr(\mathbf{x}(s)=x_{d}(s)\mathbf{e}_{d})\cdot\mathbb{E}[x_{d}(s)^{2}\mid x_{d}(s)>0]$. For $m\geq 2$, since features are independently activated with probability (1S)(1-S), a union bound over all pairs $i\neq j$ gives:
$\Pr(\|\mathbf{x}(s)\|_{0}\geq 2)\leq\sum_{i,j}\Pr(x_{i}>0,\,x_{j}>0)\leq n^{2}(1-S)^{2}.$ (35)
Since the reconstruction error is bounded by $C$, combining all cases yields:
$\left|\mathcal{L}_{\text{SDL}}-\tilde{\mathcal{L}}_{\text{SDL}}\right|\leq n^{2}C(1-S)^{2}.$ (36)
@@
Preliminary observation. By definition of $\Omega_{\mathcal{P}}$ (Definition 3.1), for any $W_{E}\in\Omega_{\mathcal{P}}$, the activation $z_{d}(W_{E}):=\sigma(W_{E}w_{p}^{d})$ satisfies:
$z_{d}(W_{E})_{i}=\begin{cases}\langle w_{E}^{i},w_{p}^{d}\rangle&\text{if }i\in\mathcal{F}_{d}\\ 0&\text{otherwise,}\end{cases}$ (37)
where $\mathcal{F}_{d}=\{i\in[n_{q}]:(\sigma(W_{E}w_{p}^{d}))_{i}>0\}$. Hence $z_{d}(W_{E})$ is linear in $W_{E}$ within $\Omega_{\mathcal{P}}$.
Per-feature decomposition. Define $f_{d}(W_{D},W_{E}):=\|w_{r}^{d}-W_{D}z_{d}(W_{E})\|^{2}$. Then:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})=\sum_{d=1}^{n}M_{d}f_{d}(W_{D},W_{E}),\quad M_{d}>0.$ (38)
Since positive linear combinations preserve biconvexity, it suffices to show each $f_{d}$ is biconvex.
Convexity in $W_{D}$. For fixed $W_{E}\in\Omega_{\mathcal{P}}$, $z_{d}$ is a fixed vector, so $f_{d}(W_{D})=\|w_{r}^{d}-W_{D}z_{d}\|^{2}$ is a squared affine function of $W_{D}$, hence convex. Its Hessian $\nabla^{2}_{W_{D}}f_{d}=2I_{n_{r}}\otimes(z_{d}z_{d}^{\top})\succeq 0$ confirms this.
Convexity in $W_{E}$. For fixed $W_{D}$, linearity of $z_{d}(W_{E})$ in $W_{E}$ gives:
$W_{D}z_{d}(W_{E})=\sum_{i\in\mathcal{F}_{d}}\langle w_{E}^{i},w_{p}^{d}\rangle w_{D}^{i},$ (39)
which is affine in $W_{E}$. Hence $f_{d}(W_{E})=\|w_{r}^{d}-W_{D}z_{d}(W_{E})\|^{2}$ is a squared affine function of $W_{E}$, which is convex. For each row $w_{E}^{i}$ with $i\in\mathcal{F}_{d}$, the Hessian $\nabla^{2}_{w_{E}^{i}}f_{d}=2\|w_{D}^{i}\|^{2}(w_{p}^{d}(w_{p}^{d})^{\top})\succeq 0$, and rows $i\notin\mathcal{F}_{d}$ do not affect $f_{d}$.
Conclusion. Each $f_{d}$ is biconvex, so $\tilde{\mathcal{L}}_{\text{SDL}}$ is biconvex over $\mathbb{R}^{n_{r}\times n_{q}}\times\Omega_{\mathcal{P}}$. ∎
@@
We analyze $\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})$ by examining the reconstruction error for each feature $d\in[n]$ active in isolation.
The encoder output is $W_{E}^{*}w_{p}^{d}=W_{p}^{\top}w_{p}^{d}$ with $k$-th component $(W_{E}^{*}w_{p}^{d})_{k}=\langle w_{p}^{k},w_{p}^{d}\rangle$. By the unit-norm condition (Assumption 2.1), $(W_{E}^{*}w_{p}^{d})_{d}=1$. Define the set of neurons activated by feature $d$:
$\mathcal{A}_{d}=\{k\in[n]:\sigma(W_{E}^{*}w_{p}^{d})_{k}>0\}.$ (40)
Since $\sigma(z)_{i}\in\{0,z_{i}\}$, we have $\sigma(W_{E}^{*}w_{p}^{d})_{k}>0$ if and only if $\langle w_{p}^{k},w_{p}^{d}\rangle>0$. Therefore $d\in\mathcal{A}_{d}$, and for $k\neq d$:
$\mathcal{A}_{d}\setminus\{d\}\subseteq\{k\in[n]:0<\langle w_{p}^{k},w_{p}^{d}\rangle\leq M\}.$ (41)
The reconstruction is $W_{D}^{*}\sigma(W_{E}^{*}w_{p}^{d})=\sum_{k\in\mathcal{A}_{d}}\langle w_{p}^{k},w_{p}^{d}\rangle w_{r}^{k}$, so the reconstruction error for feature $d$ is:
$w_{r}^{d}-W_{D}^{*}\sigma(W_{E}^{*}w_{p}^{d})$ $=w_{r}^{d}-\sum_{k\in\mathcal{A}_{d}}\langle w_{p}^{k},w_{p}^{d}\rangle w_{r}^{k}=-\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}\langle w_{p}^{k},w_{p}^{d}\rangle w_{r}^{k}.$ (42)
Applying the triangle inequality and using $\|w_{r}^{k}\|_{2}=1$ and $\langle w_{p}^{k},w_{p}^{d}\rangle\leq M$ for $k\in\mathcal{A}_{d}\setminus\{d\}$:
$\left\|w_{r}^{d}-W_{D}^{*}\sigma(W_{E}^{*}w_{p}^{d})\right\|_{2}\leq\sum_{k\in\mathcal{A}_{d}\setminus\{d\}}\langle w_{p}^{k},w_{p}^{d}\rangle\leq M(K_{d}-1),$ (43)
where $K_{d}=|\mathcal{A}_{d}|\leq n$. Squaring and summing over all features:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})\leq M^{2}\sum_{d=1}^{n}M_{d}(K_{d}-1)^{2}\leq n^{2}M^{2}\sum_{d=1}^{n}M_{d},$ (44)
where the second inequality uses $(K_{d}-1)^{2}\leq(n-1)^{2}\leq n^{2}$. ∎